# Multitask-based Evaluation of Open-Source LLM on Software Vulnerability

### v2

@@ 5 Threats to Validate
Threats to Construct Validity mainly correspond to the performance metrics in our evaluations. To minimize such threats, we consider a few widely used performance metrics to evaluate the performance of LLMs on different types of tasks. (e.g., Accuracy, Precision, Recall, and ROUGE).
@@ 6 Conclusion
This paper aims to comprehensively investigate the capabilities of LLMs for software vulnerability tasks as well as its impacts. To achieve that, we adopt a large-scale vulnerability dataset (named Big-Vul) and then conduct several experiments focusing on five dimensions: (1) Vulnerability Detection, (2) Vulnerability Assessment, (3) Vulnerability Localization, and (4) Vulnerability Description. Overall, although LLMs shows some ability in certain areas, it still needs further improvement to be competent in software vulnerability related tasks. Our research conducts a comprehensive survey of LLMs’ capabilities and provides a reference for enhancing its understanding of software vulnerabilities in the future.

### v3

@@ 5.3 Threats to Validity
Threats to Construct Validity mainly correspond to the performance metrics in our evaluations. To minimize such threats, we consider a few widely used performance metrics to evaluate the performance of LLMs on different types of tasks, e.g., Recall, Precision, and ROUGE.
@@ 6 Conclusion
This paper aims to comprehensively investigate the capabilities of LLMs for software vulnerability tasks as well as its impacts. To achieve that, we adopt a large-scale vulnerability dataset (named Big-Vul) and then conduct several experiments focusing on four dimensions: (1) Vulnerability Detection, (2) Vulnerability Assessment, (3) Vulnerability Location, and (4) Vulnerability Description. Overall, although LLMs show some ability in certain areas, they still need further improvement to be competent in software vulnerability-related tasks. Our research conducts a comprehensive survey of LLMs’ capabilities and provides a reference for enhancing its understanding of software vulnerabilities in the future.